Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SCS SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 418     Entry time: 26 Oct 2023, 15:03
Author: Sergii, Teguh 
Group: SCS 
Subject: Morning shift summary, 26.10.2023 

Morning shift summary, 26.10.2023

Achievements:

  1. Continued measuring tr-RIXS at different laser fluences

Problems:

  1. The max-jhub notebook server was very slow this morning around 9-10 h. DOC was informed, but the slowness has improved without any action from either them or our side. The reason for this slowness is still unknown, but the max-jhub server has come back to normal for now.
ELOG V3.1.4-7c3fd00